If "2+ squareroot (3)" is a root of the equation x^2 + px + q = 0 , then write the values of p and q .


Please give detailed steps.


Thanks in advance.

Grade:12

4 Answers

vikas askiitian expert
12 years ago

let A,B are the roots of this equation then

A+B = -p      ,    AB = q

if one root is 2+31/2 then other can be 2-31/2 or 31/2-2

 case 1) (A,B) = (2+roo3 , 2-root3)                                 case2) (A,B) = (root3-2) , (root3+2)

          A+B = 4 = -P                                                                       A+B = 2root3 = -p

      AB = 4-3 = 1 = q                                                                      AB = -1   = q

now we have two odered pairs for (p,q)

p,q ={ (-4,1) , (-2root3,-1) }

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

Sum of the roots=-p

Product =q

Now one of the roots is given 2+root3

You can easily finc the value of p and q

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

aryan kashyap
12 years ago

if p n q are real no.s then if 2+sqrt3 is root then 2-sqrt3 has to b d root as well.

thr4 p=-4    q=1

Sudheesh Singanamalla
12 years ago

Thanks a lot guys ! :)